OpenStudy (anonymous):

Jose's math final is worth 100 points and consists of 45 test questions. Multiple choice questions are worth 2 points and application questions are worth 4 points. How many application problems are on the test? A. 3 B. 5 C. 7 D. 9

OpenStudy (anonymous):

m + a = 45 and 2m + 4a = 100 You now just simply solve those 2 simultaneous equations. Can you do that or would you like further help?

OpenStudy (anonymous):

If you are going to attempt solving those and don't know where to start, I would suggest the elimination method. Multiply both sides of the first equation by -2 and add that new equation to the second equation. Add left side to left side and right to right.

OpenStudy (anonymous):

The "m" variable will drop out and you will be left with one equation in one variable which is: "a"

OpenStudy (anonymous):

All good now, @Reganbaughman417 ?

OpenStudy (anonymous):

Yes thank you @tcarroll010

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Reganbaughman417

OpenStudy (anonymous):

Thank you and your welcome! (:

OpenStudy (anonymous):

First let's put this in equation form. Let M stand for Multiple choice and A stand for application questions. So there are 45 questions on this test so this is the equation we have for this question. M+A=45 And We also know that the Multiple choice questions are worth 2 points and the Application 4 points and that the test is worth 100 points. So that gives us the equations: 2(M) + 4(A)=100 Now we go back to the first equations and let's subtract the application questions from 45. M+A=45 -A -A M=45-A Now we go to the second equation and substitute M in: 2(45-A) +4(A)=100 So now you distribute and solve 90-2A+4A=100 -90 -90 -2A+4A=10 2A=10 A=5 So 5 application questions